# Analytics, have some humility: a statistical view of fourth-down decision making

### v4

@@ C.1 Modeling punt outcome
Punter quality. We define a punter’s quality by a weighted sum of his punt yards over expected over all his previous punts in his career. To begin, we fit a simple expected next yardline after punting model ${{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}$ on a held-out dataset of all punts from 1999 to 2005 to avoid data bleed. We fit ${{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}$ using linear regression as a function of yardline (specifically, a cubic polynomial in yardline). Then, we define the punt yards over expected (${\mathsf{PYOE}}$) of the $n^{th}$ punt by
@@
where ${\mathsf{pq}}_{0}:=0$ and ${\mathsf{PYOE}}_{0}:=0$. We use an exponential decay weight $\alpha=0.995$ to upweight more recent punts. Finally, we standardize the punter quality column to have mean zero and standard deviation $1/2$. In Figure 7a we plot the career mean punter quality of each punter with over 250 punt attempts from 2006 to 2021. As expected, Johnny Hekker has the highest punter quality.
Punt outcome model. We use linear regression to model the expected next yardline after a punt as a function of yardline and punter quality (${\mathsf{pq}}$),
$\begin{split}{\mathbb{E}}_{\text{punt}}[\text{next yardline}]&=\ \vec{\alpha}\cdot\text{spline}\big{(}\text{yardline},\ \text{df}=4\big{)}+\ \beta_{1}\cdot{\mathsf{pq}}+\beta_{2}\cdot{\mathsf{pq}}\cdot\text{yardline}.\end{split}$ (23)
The model is trained on a dataset of $36,493$ punts from 2006 to 2021, all beyond the 30 yardline. In Figure 7b we plot the expected next yardline after a punt according to our model as a function of yardline for various punter quality values.
Figure 7: On the left, the career mean punter quality ($x$-axis) of each punter ($y$-axis) with over 250 punts from 2006 to 2021. On the right, the expected next yardline after a punt ($y$-axis) according to our model as a function of yardline ($x$-axis) for various punter quality values (color).
C.2 Modeling field goal success probability
In this Section we model field goal success probability as a function of yardline and kicker quality. It is important to adjust for kicker quality to avoid selection bias, as good kickers attempt more field goals from long distance than bad kickers. Our kicker quality adjustment, detailed below, is similar to our punter quality adjustment from the previous Section.
Kicker quality. We define a kicker’s quality by a weighted sum of his field goal probability added over all his previous kicks in his career. To begin, we fit a simple field goal probability model ${P_{\mathsf{FG}}^{(0)}}$ on a held-out dataset of all field goals from 1999 to 2005 to avoid data bleed. We fit ${P_{\mathsf{FG}}^{(0)}}$ using logistic regression as a function of yardline (specifically, a cubic polynomial spline with five degrees of freedom on the yardline). Then, we define the field goal probability added (${\mathsf{FGPA}}$) of the $n^{th}$ field goal by
@@ C.2 Modeling field goal success probability
where ${\mathsf{kq}}_{0}:=0$ and ${\mathsf{FGPA}}_{0}:=0$. As before, we use an exponential decay weight $\alpha=0.995$ to upweight more recent kicks. Finally, we standardize the kicker quality column to have mean zero and standard deviation $1/2$. In Figure 8a we plot the career mean kicker quality of each kicker with over 100 field goal attempts from 2006 to 2021. As expected, Justin Tucker has by far the highest kicker quality.
Field goal success probability model. We use logistic regression to model the probability that a kicker makes a field goal as a function of yardline and kicker quality (${\mathsf{kq}}$). Formally, our best field goal probability model is
$\begin{split}\log\bigg{(}\frac{{\mathbb{P}}(\text{make }{\mathsf{FG}})}{1-{\mathbb{P}}(\text{make }{\mathsf{FG}})}\bigg{)}=\ \vec{\alpha}\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=5\big{)}+\beta\cdot{\mathsf{kq}}.\end{split}$ (26)
Fitting this model on our dataset of $15,472$ observed field goals from 2006 to 2021 yields nontrivial probability predictions for extremely long field goals that have never before been made (e.g., nontrivial probability for a 73-yard field goal from the 55 yardline). To shrink these field goal probability predictions to zero, we impute 500 synthetic missed field goals, randomly distributed from the 51 to the 99 yardline, into our dataset. In Figure 8b we plot the probability of making a field goal according to our model as a function of yardline for various kicker quality values.
Figure 8: On the left, the career mean kicker quality ($x$-axis) of each kicker ($y$-axis) with over 100 field goal attempts from 2006 to 2021. On the right, the probability of making a field goal ($y$-axis) according to our model as a function of yardline ($x$-axis) for various kicker quality values (color).
C.3 Modeling conversion probability
In this Section we model conversion probability as a function of yards to go and team quality. First, we detail our team quality measures, which produce a more predictive conversion probability model than point spread. Then, we detail our conversion probability model.
@@ C.3 Modeling conversion probability
To avoid data bleed, we remove all plays from 1999 to 2005 from our play-by-play dataset and fit an expected points model ${\mathsf{EP^{(0)}}}$ to this held-out data.
The simple ${\mathsf{EP^{(0)}}}$ model which we fit on held-out data from 1999-2005 in order to craft our team quality metrics is a linear regression in which the outcome variable is the observed net points of the next score, a real number in $\{7,3,2,0,-2,-3,-7\}$. We model expected net points as a linear function of yardline, log yards to go, and one combined indicator for third and fourth down.
@@
where $q_{0}:=0$ and ${\mathsf{EPA^{(0)}}}_{0}:=0$. We use an exponential decay weight $\alpha$ to upweight more recent plays. We set $\alpha=0.995$ for each of our team and player quality metrics. For instance, a play which occurred 138 plays ago is weighted half as much as the previous play since $0.50.995^{138}=$ 0.5. Also, a play which occurred 459 plays ago is weighted one-tenth as much as the previous play since $0.10.995^{459}=$ 0.1. Additionally, at the start of each season, we shrink by a multiplicative factor $\gamma$. So, the shrinkage weight $\gamma_{n}$ of play $n$ is $\gamma$ if this is Mahomes’ first play of the season, otherwise it is 1. We use different shrinkage values $\gamma$ for each team and player quality metric, shown in Table 4. We tuned these $\gamma$ values to optimize predictive performance on a held-out validation set. A smaller $\gamma$ reflects a quantity that is noisier from year to year. Finally, we standardize each offensive and defensive quality metric to have mean zero and standard deviation $1/2$. In Figure 9 we plot the career mean quarterback quality of each quarterback with over 1750 attempts and whose rookie season came after 2006. As expected, Patrick Mahomes has by far the highest quarterback quality. We construct the other team quality metrics described in Table 4 in a similar fashion, via Formula (28).
@@
Table 4: Summary of team quality metrics computed via Formula (28) and their start-of-season shrinkage parameters $\gamma$. Each of these metrics apply to both the offensive and defensive teams in a play, yielding eight total metrics.
Due to a small sample size of fourth-down conversion attempts, some existing models use third down as a proxy for fourth down, as they are also high-pressure situations in which the offensive team attempts to gain a first down on that play (Romer,, 2006). There may be, however, a fundamental difference in conversion probability between third and fourth-down plays, perhaps due to psychological reasons or a different distribution of play calls on fourth down. Therefore, in our model selection process, we test models on a random 50% of fourth-down plays, and we train some models on a dataset consisting entirely of fourth-down plays and other models on a dataset consisting of third and fourth-down plays. We find that the parameters of our best conversion probability model borrow strength from third-down plays.
Our best logistic regression model adjusts for yards to go, down (third vs. fourth down), and our offensive and defensive quality metrics: quarterback quality of the offensive team (${\mathsf{qbqot}}$), non-quarterback offensive quality of the offensive team (${\mathsf{oqrot}}$), defensive quality of the defensive team against the pass (${\mathsf{dqdtap}}$), and defensive quality of the defensive team against the run (${\mathsf{dqdtar}}$). Formally, our best conversion probability model is
$\begin{split}\log\bigg{(}\frac{{\mathbb{P}}({\text{convert}})}{1-{\mathbb{P}}({\text{convert}})}\bigg{)}=&\ \vec{\alpha_{1}}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{2}}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \beta_{0}+\beta_{1}\cdot{\mathsf{qbqot}}+\beta_{2}\cdot{\mathsf{oqrot}}+\beta_{3}\cdot{\mathsf{dqdtap}}+\beta_{4}\cdot{\mathsf{dqdtar}}.\end{split}$ (29)
In Figure 10 we visualize conversion probability as a function of yards to go. We see a large spike in conversion probability with one yard to go. In Figure 11 we plot conversion probability as a function of yards to go for various values of team quality. We find that quarterback quality significantly impacts conversion probability and the other team quality measures have a small impact.
Figure 10: Conversion probability ($y$-axis) according to our model as a function of yards to go ($x$-axis), assuming average team quality.
Figure 11: Conversion probability ($y$-axis) according to our model as a function of yards to go ($x$-axis) for various values of team quality (color). Quarterback quality has by far the largest impact on conversion probability, visualized by a larger spread amongst the colored lines.
A more elaborate conversion probability model may adjust for yardline. In particular, it is plausible that it is more difficult to convert near each endzone, where space is more constricted. Additionally, a more fine-grained model would be continuous in yards to go rather than treating it as an integer. Of course, our model can only be as good as available data, which treats yards to go as an integer. But, anecdotally, fourth down and inches has a significantly higher conversion probability than fourth down and one yard to go. Finally, conversion probability depends on the offensive and defensive play call. On this view, in practice it may be better for teams to input their own conversion probability estimates as they are more informed on their play calling tendencies.
Appendix D First-down win probability model details
In our fourth-down decision procedure, we use a win probability ${\mathsf{XGBoost}}$ model fit from first-down plays. This model predicts binary win/loss as a function of score differential, game seconds remaining, point spread, yardline, receive $2^{nd}$ half kickoff indicator, offensive team’s number of timeouts remaining, defensive team’s number of timeouts remaining, total score, and
$\texttt{scoreTimeRatio}=\frac{\texttt{score\_differential}}{0.01+\texttt{game\_seconds\_remaining}}.$ (30)
Now we compare the out-of-sample predictive performance of various ${\mathsf{WP}}$ models. Our full dataset consists of all football plays from 2006 to 2021. The dataset is clustered into games, as plays from each game share the same winning team. To keep the clustered nature of our dataset intact and to avoid data bleed, we split our dataset in half by randomly sampling 50% of all games. The first-down plays from the first 50% of these games form the hold-out test set. We test on first-down plays because fourth-down decision making relies on first-down win probability. The plays from the other 50% of these games form the training set. To tune ${\mathsf{XGBoost}}$ models, we split the training set in half by randomly sampling 50% of the games from the training set. The plays from the first 50% of these games form the ${\mathsf{XGBoost}}$ training set, and the remaining plays form the validation set for hyper-parameter tuning. We then tune our ${\mathsf{XGBoost}}$ models in a similar fashion as Baldwin, (2021). We visualize the results of our model comparison in Table 5.
model
name
model
type
plays
in dataset
out-of-sample
${\mathsf{logloss}}$
Ours ${\mathsf{XGBoost}}$ first downs 0.440
Lock and Nettleton, (2014) ${\mathsf{RF}}$ all downs 0.447
Baldwin, (2021) ${\mathsf{XGBoost}}$ all downs 0.476
Table 5: Predictive performance of various ${\mathsf{WP}}$ models.
Appendix E Simulation study details
Generating plays. Formally, the outcome of the $n^{th}$ play of the $g^{th}$ game is
$\xi_{gn}\overset{iid}{\sim}\pm 1.$ (31)
The game starts at midfield, $X_{g0}=L/2$, and the game begins tied, $S_{g0}=0$. The field position at the start of play $n$ is
$X_{g,n+1}:=\begin{cases}X_{gn}+\xi_{gn}&\text{ if }0<X_{gn}+\xi_{gn}<L\text{ (not a TD)}\\ L/2&\text{ else,}\end{cases}$ (32)
and the score differential at the start of play $n$ is
$S_{g,n+1}:=\begin{cases}S_{gn}+1&\text{ if }X_{gn}+\xi_{gn}=0\text{ (TD)}\\ S_{gn}-1&\text{ if }X_{gn}+\xi_{gn}=L\text{ (opp. TD)}\\ S_{gn}&\text{ else.}\end{cases}$ (33)
The response column win is
$y_{gn}\equiv y_{g,N+1}:=\begin{cases}1&\text{ if }S_{g,N+1}>0\\ 0&\text{ if }S_{g,N+1}<0\\ \text{Bernoulli}(1/2)&\text{ else (overtime). }\end{cases}$ (34)
As in our dataset of real football plays, this response column is highly correlated – plays from the same game share the same draw of the winner of the game.
Generating observational data. We create a dataset of plays from $G$ games. Each game consists of $N$ plays, and the field consists of $L$ yardlines. The results from each game yield a simulated dataset
${\mathscr{D}}=\{(n,X_{gn},S_{gn},y_{gn}):n=1,...,N\text{ and }g=1,...,G\}.$ (35)
True win probability. The true win probability
${\mathsf{WP}}(n,x,s):={\mathbb{P}}(S_{g,N+1}>0|X_{gn}=x,S_{gn}=s)$ (36)
of our simplified version of football is computed explicitly using dynamic programming,
${\mathsf{WP}}(N+1,x,s)=\begin{cases}1&\text{if }s>0\\ 1/2&\text{if }s=0\\ 0&\text{if }s<0,\\ \end{cases}$ (37)
and
${\mathsf{WP}}(n-1,x,s)=\begin{cases}\frac{1}{2}{\mathsf{WP}}(n,\frac{L}{2},s+1)+\frac{1}{2}{\mathsf{WP}}(n,x+1,s)&\text{if }x=1\\ \frac{1}{2}{\mathsf{WP}}(n,x-1,s)+\frac{1}{2}{\mathsf{WP}}(n,\frac{L}{2},s-1)&\text{if }x=L-1\\ \frac{1}{2}{\mathsf{WP}}(n,x-1,s)+\frac{1}{2}{\mathsf{WP}}(n,x+1,s)&\text{else.}\end{cases}$ (38)
Appendix F Fitting our fourth-down decision procedure
In this Section we describe in detail our fourth-down decision procedure, which incorporates uncertainty quantification.
We begin by fitting the traditional fourth-down decision procedure. We fit our first-down win probability model, an ${\mathsf{XGBoost}}$ model fit from first-down plays, to our dataset (recall Section 2.4 and Appendix D). We also fit our decision transition models from our dataset (recall the punt outcome model, field goal success probability model, and conversion probability model, described in Appendix C). We then knit these models together to estimate fourth-down win probability for each decision ${\mathsf{Go}}$, ${\mathsf{FG}}$, and ${\mathsf{Punt}}$ (described in Section 2 and Appendix B). The estimated optimal decision is the one that maximizes win probability.
We use bootstrapping to quantify uncertainty in our fourth-down decision recommendations. We generate $B=100$ bootstrapped datasets using a randomized cluster bootstrap as follows (we justify our choice of $B=100$ in Appendix G). Let $G$ denote the number of games in our observed play-by-play dataset. To generate a bootstrapped dataset we re-sample $G\cdot\phi$ games uniformly with replacement. We use $\phi=0.5$ as justified by the simulation study from Section 3.1 in order to achieve adequate marginal coverage. Within each game $g$, which contains $N_{g}$ drives, we re-sample $N_{g}$ drives uniformly with replacement. We re-sample drives with replacement, rather than plays with replacement as we did in the simulation study, since it provides a more realistic re-sampling of real football data.
From each of the $B$ bootstrapped datasets we fit a first-down win probability model. We knit each bootstrapped first-down win probability model together with the decision transition models fit from observed data to form $B$ bootstrapped fourth-down decision models. For each game-state, from the $B$ bagged decisions we calculate bootstrap percentage and form a $90\%$ confidence interval on the estimated gain in win probability.
We just bootstrap the first-down win probability model and not the decision transition models because the former model is subject to much higher uncertainty. Decision transition probability models are fit from i.i.d. play-level outcomes, whereas first-down win probability models are fit from clustered, highly correlated game-level outcomes. On this view, as discussed in Section 2.3, estimating first-down win probability is the most difficult component of fourth-down decision making and is the source of the majority of uncertainty in fourth-down win probability estimates. Hence in this paper we focus on propagating uncertainty from first-down win probability models into the fourth-down decision procedure.
Appendix G Tuning the number of bootstrapped datasets
As discussed in Section 3.2, we use bootstrap percentage to incorporate uncertainty quantification into fourth-down decision making because it quantifies uncertainty in the decision itself. Although the strength of a decision is proportional to the ${\mathsf{boot\%}}$, we cannot necessarily rely on a non-technical football coach to fully process how ${\mathsf{boot\%}}$ maps to the strength of a decision recommendation. To facilitate easier communication to a coach, we bin decision recommendations into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$).
@@ Appendix G Tuning the number of bootstrapped datasets
Since we ultimately bucket decision recommendations into three bins (see Section 4.2), we want the number of bootstrapped models $B$ to be large enough to stably categorize decisions into one of these three bins. We also want $B$ to be as small as possible in order to quickly evaluate a fourth-down recommendation during a football game. Hence, we conduct a stability analysis to choose $B$.
@@
For each category, $p_{i}^{\text{category, }(B)}$ is the proportion of the $M$ bootstrap draws that play $i$ is put in that category. For our procedure with $B$ bootstrapped models to be stable, we want $p_{i}^{\text{category, }(B)}$ to be close to 1 for each category, which means that a play’s categorization is not dependent on the randomness inherent in generating $B$ bootstrapped models. Hence, we want $p_{i}^{(B)}$, the maximum $p_{i}^{\text{category, }(B)}$ across the three categories, to be close to 1. Hence, we want the mean across all plays $\overline{p}^{(B)}=\frac{1}{n}\sum_{i=1}^{n}p_{i}^{(B)}$ to be close to 1. In Figure 12 we see that for $B=100$, $\overline{p}^{(B)}=0.9$ and the vast majority of plays $i$ have $p_{i}^{(B)}=1$. Some plays have $p_{i}^{(B)}$ lower than 1 since they lie near the border of two categories. We believe $\overline{p}^{(B)}=0.9$ is sufficiently large and hence use $B=100$ in our fourth-down decision procedure.
Figure 12: This is a histogram of $\{p_{i}^{(B)}\}_{i}$. The blue line is the mean. Since most of the value equals 1 and the mean is high enough, $B=100$ is large enough.
Appendix H Baseline coaches’ decision model
To compare our decision-making procedure to the decisions that actual football coaches tend to make, we model the probability that a coach chooses a decision in $\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ as a function of game-state. We use ${\mathsf{XGBoost}}$ to fit these coach probabilities. ${\mathsf{XGBoost}}$ works well here because we have $94,786$ fourth-down plays in our full dataset of plays since 1999, and each play is an independent observation of a coach’s decision. In particular, we fit these coach probabilities as a function of yardline, yards to go, game seconds remaining, score differential, point spread, and era (1999-2001, 2002-2005, 2006-2013, 2014-2017, and 2018-present). In Figure 13 we visualize these coach decision models, and the results make intuitive sense. For the most part, coaches punt deep in their own territory and kick field goals near the opponent’s endzone, except for with one and sometimes two yards to go. Also, at the end of the game, coaches’ decision-making changes depending on the number of points they need to score to win the game.

### v5

@@ Appendix B Estimating player / team quality
Punter quality. We define a punter’s quality by a weighted sum of his punt yards over expected over all his previous punts in his career. To begin, we fit a simple expected next yardline after punting model ${{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}$ on a held-out dataset of all punts from 1999 to 2005. We fit ${{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}$ using linear regression as a function of yardline (specifically, a cubic polynomial in yardline). Then, we define the punt yards over expected (${\mathsf{PYOE}}$) of the $n^{th}$ punt by
@@
where ${\mathsf{pq}}_{0}:=0$ and ${\mathsf{PYOE}}_{0}:=0$. We use an exponential decay weight $\alpha=0.995$ to upweight more recent punts. For instance, a play which occurred 138 plays ago is weighted half as much as the previous play since $0.50.995^{138}=$ 0.5. Also, a play which occurred 459 plays ago is weighted one-tenth as much as the previous play since $0.10.995^{459}=$ 0.1. Finally, we standardize the punter quality column to have mean zero and standard deviation $1/2$. In Figure 9a we plot the career mean punter quality of each punter with over 250 punt attempts from 2006 to 2021. As expected, Johnny Hekker has the highest punter quality.
Figure 9: (a) The career mean punter quality ($x$-axis) of each punter ($y$-axis) with over 250 punts from 2006 to 2021. (b) The career mean kicker quality ($x$-axis) of each kicker ($y$-axis) with over 100 field goal attempts from 2006 to 2021.
Kicker quality. We define a kicker’s quality by a weighted sum of his field goal probability added over all his previous kicks in his career. To begin, we fit a simple field goal probability model ${P_{\mathsf{FG}}^{(0)}}$ on a held-out dataset of all field goals from 1999 to 2005. We fit ${P_{\mathsf{FG}}^{(0)}}$ using logistic regression as a function of yardline (specifically, a cubic spline with five degrees of freedom on the yardline). Then, we define the field goal probability added (${\mathsf{FGPA}}$) of the $n^{th}$ field goal by
@@
where ${\mathsf{kq}}_{0}:=0$ and ${\mathsf{FGPA}}_{0}:=0$. As before, we use an exponential decay weight $\alpha=0.995$ to upweight more recent kicks. Finally, we standardize the kicker quality column to have mean zero and standard deviation $1/2$. In Figure 9b we plot the career mean kicker quality of each kicker with over 100 field goal attempts from 2006 to 2021. As expected, Justin Tucker has by far the highest kicker quality.
@@
We remove all plays from 1999 to 2005 from our play-by-play dataset and fit an expected points model ${\mathsf{EP^{(0)}}}$ to this held-out data. In fitting this model, the outcome variable is the observed net points of the next score, a real number in $\{7,3,2,0,-2,-3,-7\}$. We model expected net points as a linear function of yardline, log yards to go, and one combined indicator for third and fourth down.
@@
where $q_{0}:=0$ and ${\mathsf{EPA^{(0)}}}_{0}:=0$. We use an exponential decay weight $\alpha$ to upweight more recent plays. We set $\alpha=0.995$ for each of our team and player quality metrics to upweight more recent plays. Additionally, at the start of each season, we shrink by a multiplicative factor $\gamma$. So, the shrinkage weight $\gamma_{n}$ of play $n$ is $\gamma$ if this is Mahomes’ first play of the season, otherwise it is 1. We use different shrinkage values $\gamma$ for each team and player quality metric, shown in Table 3. We tuned these $\gamma$ values to optimize predictive performance on a held-out validation set. A smaller $\gamma$ reflects a quantity that is noisier from year to year. Finally, we standardize each offensive and defensive quality metric to have mean zero and standard deviation $1/2$. In Figure 10 we plot the career mean quarterback quality of each quarterback with over 1750 attempts and whose rookie season came after 2006. As expected, Patrick Mahomes has by far the highest quarterback quality. We construct the other team quality metrics described in Table 3 in a similar fashion, via Formula (26).
@@
Table 3: Summary of team quality metrics computed via Formula (26) and their start-of-season shrinkage parameters $\gamma$.
Appendix C Tuning the number of bootstrapped datasets
@@
We use ${\mathsf{boot\%}}$ to measure uncertainty in the estimated optimal decision, but we cannot necessarily rely on a non-technical football coach to fully process how ${\mathsf{boot\%}}$ maps to the strength of a recommendation. To facilitate easier communication to a coach, we bin decision recommendations into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$). Since we ultimately bucket decision recommendations into these three bins (see Section 4.2), we want the number of bootstrapped models $B$ to be large enough to stably categorize decisions into one of these three bins. We also want $B$ to be as small as possible in order to quickly evaluate a fourth-down recommendation during a football game. Hence, we conduct a stability analysis to choose $B$.
@@
For each category, $p_{i}^{\text{category, }(B)}$ is the proportion of the $M$ bootstrap draws that play $i$ is put in that category. For our procedure with $B$ bootstrapped models to be stable, we want $p_{i}^{\text{category, }(B)}$ to be close to 1 for each category, which means that a play’s categorization is not dependent on the randomness inherent in generating $B$ bootstrapped models. Hence, we want $p_{i}^{(B)}$, the maximum $p_{i}^{\text{category, }(B)}$ across the three categories, to be close to 1. Hence, we want the mean across all plays $\overline{p}^{(B)}=\frac{1}{n}\sum_{i=1}^{n}p_{i}^{(B)}$ to be close to 1. In Figure 11 we see that for $B=101$, $\overline{p}^{(B)}=0.9$ and the vast majority of plays $i$ have $p_{i}^{(B)}=1$. Some plays have $p_{i}^{(B)}$ lower than 1 since they lie near the border of two categories. We believe $\overline{p}^{(B)}=0.9$ is sufficiently large and hence use $B=101$ in our fourth-down decision procedure.
Figure 11: This is a histogram of $\{p_{i}^{(B)}\}_{i}$. The blue line is the mean. Since most of the values equal 1 and the mean is high enough, $B=101$ is large enough.
Appendix D Baseline coaches’ decision model
To compare our decision-making procedure to the decisions that actual football coaches tend to make, we model the probability that a coach chooses a decision in $\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ as a function of game-state. We use ${\mathsf{XGBoost}}$ to fit these coach probabilities. ${\mathsf{XGBoost}}$ works well here because we have $94,786$ fourth-down plays in our full dataset of plays since 1999, and each play is a reasonably approximate independent observation of a coach’s decision. In particular, we fit these coach probabilities as a function of yardline, yards to go, game seconds remaining, score differential, point spread, and era (1999-2001, 2002-2005, 2006-2013, 2014-2017, and 2018-present). In Figure 12 we visualize these coach decision models, and the results make intuitive sense. For the most part, coaches punt deep in their own territory and kick field goals near the opponent’s endzone, except for with one and sometimes two yards to go. Also, at the end of the game, coaches’ decision-making changes depending on the number of points they need to score to win the game.